OpenStudy (emilyisanemoji):

Solve the equation for x 9x = -5y

OpenStudy (emilyisanemoji):

help?

OpenStudy (retireed):

How would you isolate x to one side of the equation?

OpenStudy (emilyisanemoji):

im not very good at math, and i really dont know

OpenStudy (photon336):

when you solve for x you must get x by itself

OpenStudy (emilyisanemoji):

so 9 divided by -5y

OpenStudy (retireed):

if I had.... x+4 = 7y I would isolate the x by subtracting 4 from both side like x + 4 - 4 = 7y - 4 this would equal x = 7y - 4 Now x is isolated to the left side of the equation What do YOU have to do to isolate x to the left side of your equation?? 9x = -5y

OpenStudy (emilyisanemoji):

idk

OpenStudy (retireed):

Take a guess

OpenStudy (emilyisanemoji):

divide 9 by -5, it just seems reasonable to me.

OpenStudy (retireed):

That seems reasonable to me, now write out the equation like I did above so I can see if you understand the process